# Air-Ground Collaborative Robots for Fire and Rescue Missions: Towards Mapping and Navigation Perspective

### v1

@@ Abstract
Air-ground collaborative robots have shown great potential in the field of fire and rescue, which can quickly respond to rescue needs and improve the efficiency of task execution. Mapping and navigation, as the key foundation for air-ground collaborative robots to achieve efficient task execution, have attracted a great deal of attention. This growing interest in collaborative robot mapping and navigation is conducive to improving the intelligence of fire and rescue task execution, but there has been no comprehensive investigation of this field to highlight their strengths. In this paper, we present a systematic review of the ground-to-ground cooperative robots for fire and rescue from a new perspective of mapping and navigation. First, an air-ground collaborative robots framework for fire and rescue missions based on unmanned aerial vehicle (UAV) mapping and unmanned ground vehicle (UGV) navigation is introduced. Then, the research progress of mapping and navigation under this framework is systematically summarized, including UAV mapping, UAV/UGV co-localization, and UGV navigation, with their main achievements and limitations. Based on the needs of fire and rescue missions, the collaborative robots with different numbers of UAVs and UGVs are classified, and their practicality in fire and rescue tasks is elaborated, with a focus on the discussion of their merits and demerits. In addition, the application examples of air-ground collaborative robots in various firefighting and rescue scenarios are given. Finally, this paper emphasizes the current challenges and potential research opportunities, rounding up references for practitioners and researchers willing to engage in this vibrant area of air-ground collaborative robots.
@@ I Introduction
With the development of robotics and artificial intelligence technology, air-ground collaborative robots based on unmanned aerial vehicle (UAV) and unmanned ground vehicle (UGV) have been proven to be effective in fire and rescue missions [4, 5, 6, 7]. It can provide more information support and operational flexibility for fire and rescue, which is of great significance to improving the efficiency and intelligence of search and rescue missions [8]. For instance, in fire scene investigation, UAVs can provide high-altitude perspectives and global observations to help firefighters understand the spread of the fire, the location of the fire source, and the degree of smoke, etc., while UGVs can search on the nearby ground to provide close-range fire monitoring and data collection, especially for complex terrain or areas that require detailed searches. In addition, since the operating costs of UAVs and UGVs are relatively low, the air-ground collaborative robots can effectively save the use of search and rescue resources, as well as reduce the risks and possibility of injuries to rescuers [9, 10]. To this end, air-ground collaborative robots have the advantages of UAV and UGV, and have been widely used in fire and rescue missions through mutual cooperation.
For air-ground collaborative robots, mapping and navigation technology, as the key foundation for realizing efficient and autonomous collaborative working between UAV and UGV, has received increasing amounts of attention [11, 12, 13, 14, 15]. In general, with the help of the superior mobility of UAV to quickly perceive large-scale environmental information, on the one hand, a coarse-grained full-scene map for UGV navigation is established, and on the other hand, the mission scene is efficiently obtained [9, 16]. Based on this, the UGV can achieve rapid path planning and safe navigation to perform fire and rescue missions [17, 18]. During this period, refined modeling can be achieved in combination with the local information perceived by the UGV to provide more comprehensive and multi-angle environmental information. Therefore, the collaborative working paradigm of air-ground robots is a promising solution to improve the efficiency, accuracy and safety of fire and rescue missions.
@@
Different from the above existing survey literature, this paper reviews the mapping and navigation methods of air-ground collaborative robots for fire and rescue missions from a new perspective, which is guided by the requirements of UAVs/UGVs to perform fire and rescue operations in a collaborative manner. First, a system architecture of air-ground collaborative robots for fire and rescue is introduced. Then, within this framework, we discuss the different types of maps built by UAV for UGV navigation, focusing on what kind of maps are suitable for UGV navigation. On this basis, the path planning and navigation methods of UGV are outlined from the perspective of different map types. Furthermore, for fire and rescue tasks, the air-ground collaborative robots are divided into four categories based on the number of UAVs and UGVs, and their different characteristics are elaborated. In addition, the case analysis of air-ground collaborative robot operations in fire and rescue scenarios is provided. Finally, the existing challenges and future research directions are discussed. In summary, the contributions of this paper mainly include:
@@
The remainder of this work is organized as follows. Section II introduces the air-ground collaborative navigation framework based on the requirements of performing fire and rescue missions. In Section III, the different types of maps constructed by UAVs in air-ground collaborative systems are discussed, which is the basis for UGV navigation. Section IV summarizes the navigation methods of UGVs from the perspectives of positioning and path planning, while Section V, the collaborative navigation system is classified and discussed according to the different numbers of UAVs/UGVs in air-ground collaboration. In Section VI, the application of mapping and navigation in fire and rescue scenarios is presented, and specific cases are explained. Finally, the conclusion of the paper and the trends for future research of air-ground collaborative mapping and navigation in the field of fire and rescue are given in Section VII.
@@ II System architecture of air-ground collaborative robots
Air-ground collaborative robots play an increasingly irreplaceable role in fire and rescue,which can effectively improve the efficiency, accuracy and safety of fire and rescue operations. As shown in Fig. 1, once a fire occurs, air-ground collaborative robots can quickly enter the fire scenario to perceive the location, scale and environmental conditions of the fire, and perform fire and rescue missions. According to mission requirements, environment mapping and navigation are the key foundations for achieving this goal. Typically, UAVs and need to be able to quickly build environmental maps thanks to their large-scale environmental perception capabilities, assisting AGVs in planning and navigation in challenging environments [29].
@@ III UAV Map Building for UGV Navigation
2-D Grid Map Li et al. [35] Enhancing obstacle detection via post-correction ground mapping - Digital camera Disaster-rescue scenario (Simulation)
Chatziparaschis et al. [36] Constructing a 2-D map, integrated with 3D OctoMap, to produce a navigable map for UGVs DJI Matrice 100 Hokuyo UTM-30LX-EW LiDAR, Stereo camera University campus
@@
Pei et al. [41] Achieving air-ground cooperative point cloud map fusion construction relying on LiDAR - LiDAR Simulation environment
@@
Blochliger et al. [45] Extracting occupancy information from sparse point clouds for a 3-D topological map - Visual-inertial sensor Warehouse, Office, Open space (Datasets)
Gomez et al. [46] Using the Free-area graph and the Topological map to represent information - Hokuyo URG-04LX laser sensor, Asus Xtion depth camera Simulation environment / Real-world office
Semantic Map Vasić et al. [47] Segmenting images collected by UAVs with neural networks to construct environmental maps DJI Phantom 3, DJI Mavic Pro Camera Car parks
Yue et al. [48] Constructing a semantic map through a hierarchical cooperative probabilistic semantic mapping framework UAV (with Intel NUC 6i7KYK @ 2.60 GHz CPU) ZED stereo camera Open Carpark, Indoor-Outdoor mixed environment
Bultmann et al. [49] Integrating multi-modal semantic fusion and label propagation techniques to build a semantic map in real-time DJI Matrice 210 v2 Ouster OS0-128 3D-LiDAR, Intel RealSense D455 RGB-D cameras Urban environment, Disaster test site
@@ IV-A UAV and UGV Co-Localization
Multi-sensor fusion localization Donati et al. [80] Fusing GPS, IMU, and ultrasonic sensor data with 3-D georeferenced map information Lack of specific evaluation of the overall system real-time performance
Kayhani et al. [81] Based on camera and IMU; Considering the cost-effectiveness and lightweight design of the localization method Need to manually place or replace the label in the scene
@@ IV-B UGV Navigation
2-D grid maps can be directly used for UGV navigation, allowing the UGV to plan the optimal path from the current position to the target position [103]. Representatively, Dijkstra [109] proposed a Dijkstra algorithm for shortest path, which solves the single-source shortest path problem of weighted directed or undirected graphs by using breadth-first search. In order to improve the operation speed of Dijkstra, Hart et al. [110] developed an A* algorithm, which uses a heuristic function to predict the target position, thereby giving priority to exploring areas that appear to be closer to the target. This allows the robot to find the target in fewer steps. In response to the problems of long calculation time and many search nodes when A* is applied to global path planning, a variety of improved algorithms based on search and sampling are proposed for UGV planning and navigation (refer to Fig. 9) [111, 104, 112, 113]. Although the UGV navigation method based on 2-D grid can achieve efficient and safe navigation, its navigation performance is subject to the accuracy of the map, and it has limitations such as poor safety and adaptability to 3-D obstacles and highly dynamic scenes.
@@ IV-B3 Topological map-based navigation
UGV can directly use the topological map constructed by UAV to plan the reference line for navigation, which can effectively reduce the path planning time [63, 116]. Blochliger et al. [45] proposed Topomap, a framework for creating multifunctional topological maps and reliable navigation, which uses a graph search algorithm to traverse topological points to obtain the shortest path. Wang et al. [43] used a graph search algorithm to search for a global path based on the 2-D topological map drawn by the UAV, and obtained a safe and feasible global path to achieve autonomous navigation of the UGV. Oleynikova et al. [117] constructed a sparse graph containing 3-D topological information for robot planning and navigation in a 3-D environment, as shown in Fig. 11. Ravankar et al. [105] proposed a hybrid topological mapping and navigation method for mobile robots, using the Djikstra and A* algorithms to perform navigation between topological nodes. Zuo et al. [106] introduced the Voronoi diagram into the topological mapping model. On this basis, a three-stage fast path search and navigation method was designed. Topological maps allow UGVs to perform efficient path planning and navigation. However, due to the large amount of debris in fire and rescue scenarios, it is difficult for topological maps to accurately draw the obstacles and feasible paths, making it challenging for UGVs to safely navigate based on topological maps.
@@ V-B A single UAV with multi-UGVs
The collaboration of a UAV with multi-UGVs can improve the efficiency of mission execution and provide faster response capabilities. In this configuration, the UAV can provide comprehensive environmental information to UGVs, allowing UGVs to collaboratively perform search and rescue missions [133]. In addition, UGVs can also share map data, sensor information, path planning, etc. to improve collaborative work efficiency.
Miller et al. [50] proposed a semantic-based air-ground collaborative robot system framework that can realize collaborative navigation of a UAV and multiple UGVs. The system deploys a UAV to create a semantic map of the ground in real time, which is then used for UGVs’ positioning and semantic navigation tasks. To improve the navigation efficiency of multi-UGVs missions, Zhang et al. [134] suggested a strategy to organize multi-UGVs formations. In this system, a frontier-based exploration strategy is used to search for cooperative targets on the map, and the A* algorithm is then employed to plan the collision-free path through all targets. Similarly, Zang et al. [135] designed a multi-UGVs system solution based on coordinated behavior planning and trajectory planning. However, the coordination problem between UAV and multiple UGVs is not considered. In addition, in the fire and rescue missions, UGVs may be scattered over a large area, which may affect their collaboration due to the limited perception capabilities. In this case, UAV not only needs to perform map construction to assist UGV navigation, but also needs to act as a messenger to support indirect communication of UGV. Yulong et al. [123] addressed the problem of path planning for a messenger UAV in a system involving aerial and ground collaboration with multiple UGVs. This provides a solution for achieving one-to-many air-ground collaborative navigation, facilitating real-time information sharing and coordinated rescue operations in fire and rescue scenarios. solve the path planning problem of collaborative heterogeneous robots consisting of a UAV and multiple UGVs, Ding et al. [124] developed an effective memetic algorithm to find the shortest route that allows the messenger UAV to access all moving UGVs. This treatment allows the UGVs to perform fire and rescue missions in a collective manner over a large area (refer to Fig. 14), and improve mission efficiency to a certain extent. However, for large-scale and complex fire and rescue scenarios, this kind of a single UAV with multi-UGVs collaboration requires UAV to have long-term perception and environmental modeling capabilities, which in turn makes the single UAV face challenges such as long flight time.
@@ V-C Multi-UAVs with a single UGV
The configuration of multi-UAVs with a single UGV is suitable for fire and rescue missions that require large-scale monitoring and data collection. Multiple UAVs work together to provide a more comprehensive perspective and perception capabilities, which can complete map construction faster to support UGV navigation and task execution. In addition, multiple UAVs also increase the fault tolerance and robustness of the collaborative system to a certain extent. Even if one UAV fails, other UAVs can still continue to perform the mission. This redundancy reduces the impact of UAV failures on mission execution.
Stampa et al. [125] designed an air-ground collaborative system consisting of multiple UAVs and a UGV as a mobile base station, as depicted in Fig. 15. In this system, the data collected by the UAVs are fused together, and then a 3-D map of the environment is constructed based on SLAM technology to assist the navigation of robots and human rescue missions. When the accurate environmental map is completed, the system switches to monitoring mode, with some UAVs providing a real-time aerial overview of the scene and quickly detecting environmental changes to effectively address the complexity and dynamics of the fire and rescue environment. Scaramuzza et al. [136] described a collaborative system consisting of multiple UAVs that can achieve autonomous navigation in large, unknown and GPS-free environments. These works mainly involve map construction of multi-UAV systems, but pay less attention to UGV navigation based on maps. For the path planning problem in multi-UAVs-UGV collaborative mode, Ren et al. [126] developed a framework for path planning in an air-ground collaborative system consisting of a pair of UAVs and a UGV. This architecture can be used for tasks such as collaborative search and rescue of heterogeneous robotic systems. Regarding the collaborative operation of UGV-UAV, Peng et al. [127] gave a routing and scheduling algorithm for multiple UAVs and a UGV, which allows multiple UAVs carried by the UGV to perform rescue tasks such as delivering multiple packages in different locations at the same time. The above work supports the collaborative execution of firefighting and rescue missions by UAVs and a UGV in large-scale scenarios. Although multiple UAVs can provide efficient environmental perception and map construction, the efficiency and timeliness of a single UGV in executing such emergency tasks cannot be guaranteed [137].

### v2

@@ Abstract
Air-ground collaborative robots have shown great potential in the field of fire and rescue. Mapping and navigation, as the key foundation for air-ground collaborative robots to achieve efficient task execution, have attracted a great deal of attention. This growing interest in collaborative robot mapping and navigation is conducive to improving the intelligence of fire and rescue task execution, but there has been no comprehensive investigation of this field to highlight their strengths. In this paper, we present a systematic review of the ground-to-ground cooperative robots for fire and rescue from a new perspective of mapping and navigation. First, an air-ground collaborative robots framework for fire and rescue missions based on unmanned aerial vehicle (UAV) mapping and unmanned ground vehicle (UGV) navigation is introduced. Then, the research progress of mapping and navigation under this framework is systematically summarized, including UAV mapping, UAV/UGV co-localization, and UGV navigation, with their main achievements and limitations. Based on the needs of fire and rescue missions, the collaborative robots with different numbers of UAVs and UGVs are classified, and their practicality in fire and rescue tasks is elaborated, with a focus on the discussion of their merits and demerits. In addition, the application examples of air-ground collaborative robots in various firefighting and rescue scenarios are given. Finally, this paper emphasizes the current challenges and potential research opportunities, rounding up references for practitioners and researchers willing to engage in this vibrant area of air-ground collaborative robots.
@@ I Introduction
With the development of robotics and artificial intelligence technology, air-ground collaborative robots based on unmanned aerial vehicle (UAV) and unmanned ground vehicle (UGV) have been proven to be effective in fire and rescue missions [4, 5, 6, 7]. It can provide more information support and operational flexibility for fire and rescue, which is of great significance to improving the efficiency and intelligence of search and rescue missions [8]. For instance, in fire scene investigation, UAVs can provide high-altitude perspectives and global observations to help firefighters understand the spread of the fire, and the location of the fire source, etc., while UGVs can search on the nearby ground to provide close-range fire monitoring and data collection, especially for complex terrain or areas that require detailed searches. In addition, since the operating costs of UAVs and UGVs are relatively low, the air-ground collaborative robots can effectively save the use of search and rescue resources, as well as reduce the risks and possibility of injuries to rescuers [9, 10]. To this end, air-ground collaborative robots have the advantages of UAV and UGV, and have been widely used in fire and rescue missions through mutual cooperation.
For air-ground collaborative robots, mapping and navigation technology, as the key foundation for realizing efficient and autonomous collaborative working between UAV and UGV, has received increasing amounts of attention [11, 12, 13, 14, 15]. In general, with the help of the superior mobility of UAV to quickly perceive large-scale environmental information, on the one hand, a coarse-grained full-scene map for UGV navigation is established, and on the other hand, the mission scene is efficiently obtained [16]. Based on this, the UGV can achieve rapid path planning and safe navigation to perform fire and rescue missions [8, 17]. During this period, refined modeling can be achieved in combination with the local information perceived by the UGV to provide more comprehensive environmental information. Therefore, the collaborative working paradigm of air-ground robots is a promising solution to improve the efficiency, accuracy and safety of fire and rescue missions.
@@
Different from the above existing survey literature, this paper reviews the mapping and navigation methods of air-ground collaborative robots for fire and rescue missions from a new perspective, which is guided by the requirements of UAVs/UGVs to perform fire and rescue operations in a collaborative manner. First, a system architecture of air-ground collaborative robots for fire and rescue is introduced. Then, within this framework, we discuss the different types of maps built by UAV for UGV navigation, focusing on their applicability for UGV navigation. On this basis, the path planning and navigation methods of UGV are outlined from the perspective of different map types. Furthermore, for fire and rescue tasks, the air-ground collaborative robots are divided into four categories based on the number of UAVs and UGVs, and their different characteristics are elaborated. In addition, the case analysis on the application of air-ground collaborative robots in fire and rescue scenarios is provided. Finally, the existing challenges and future research directions are discussed. The contributions of this paper mainly include:
@@
The remainder of this work is organized as follows. Section II introduces the air-ground collaborative navigation framework based on the requirements of performing fire and rescue missions. In Section III, the different types of maps constructed by UAVs for UGV navigation are discussed, which is the basisn. Section IV summarizes the navigation methods of UGVs from the perspectives of positioning and path planning, while Section V classifies and discusses the collaborative navigation system according to the different numbers of UAVs/UGVs. In Section VI, the application of mapping and navigation in fire and rescue scenarios is presented, and specific cases are explained. Finally, the conclusion of the paper and the trends for future research of air-ground collaborative mapping and navigation in the field of fire and rescue are given in Section VII.
@@ II System architecture of air-ground collaborative robots
Air-ground collaborative robots play an increasingly irreplaceable role in fire and rescue, which can effectively improve the efficiency and safety of fire and rescue operations. As shown in Fig. 1, once a fire occurs, air-ground collaborative robots can quickly enter the fire scenario to perceive the location, scale and environmental conditions of the fire, and perform fire and rescue missions. According to mission requirements, environment mapping and navigation are the key foundations for achieving this goal. Typically, UAVs and need to be able to quickly build environmental maps thanks to their large-scale environmental perception capabilities, assisting UGVs in planning and navigation in challenging environments [28].
@@ III UAV Map Building for UGV Navigation
2-D Grid Map Chatziparaschis et al. [34] Constructing a 2-D map, integrated with 3D OctoMap, to produce a navigable map for UGVs DJI Matrice 100 Hokuyo UTM-30LX-EW LiDAR, Stereo camera University campus
@@
Gomez et al. [42] Using the Free-area graph and the Topological map to represent information - Hokuyo laser sensor, Asus Xtion camera Simulation environment / Real-world office
Semantic Map Yue et al. [43] Constructing a semantic map through a hierarchical cooperative probabilistic semantic mapping framework UAV (with Intel NUC 6i7KYK @ 2.60 GHz CPU) ZED stereo camera Open Carpark, Indoor-Outdoor mixed environment
Bultmann et al. [44] Integrating multi-modal semantic fusion and label propagation techniques to build a semantic map in real-time DJI Matrice 210 v2 3D-LiDAR, RealSense D455 Urban environment, Disaster test site
@@ IV-A UAV and UGV Co-Localization
Multi-sensor fusion localization Kayhani et al. [79] Based on camera and IMU; Considering the cost-effectiveness and lightweight design of the localization method Need to manually place or replace the label in the scene
@@ IV-B UGV Navigation
2-D grid maps can be directly used for UGV navigation, allowing the UGV to plan the optimal path from the current position to the target position [102]. Representatively, Dijkstra [108] proposed a Dijkstra algorithm for shortest path, which solves the single-source shortest path problem of weighted directed or undirected graphs by using breadth-first search. In order to improve the operation speed of Dijkstra, Hart et al. [109] developed an A* algorithm, which uses a heuristic function to predict the target position, thereby giving priority to exploring areas that appear to be closer to the target. This allows the robot to find the target in fewer steps. In response to the problems of long calculation time and many search nodes in A*-based global path planning, a variety of improved algorithms based on search and sampling are proposed for UGV planning and navigation (refer to Fig. 9) [110, 103, 111, 112]. Although the UGV navigation method based on 2-D grid can achieve efficient and safe navigation, its navigation performance is subject to the accuracy of the map, and it has limitations such as poor adaptability to 3-D obstacles and highly dynamic scenes.
@@ IV-B3 Topological map-based navigation
UGV can directly use the topological map constructed by UAV to plan the reference line for navigation, which can effectively reduce the path planning time [60, 115]. Blochliger et al. [62] proposed Topomap, a framework for creating multifunctional topological maps and reliable navigation, which uses a graph search algorithm to traverse topological points to obtain the shortest path. Wang et al. [40] used a graph search algorithm to search for a global path based on the 2-D topological map drawn by the UAV, and obtained a safe and feasible global path to achieve autonomous navigation of the UGV. Oleynikova et al. [116] constructed a sparse graph containing 3-D topological information for robot planning and navigation in a 3-D environment, as shown in Fig. 11. Ravankar et al. [104] proposed a hybrid topological mapping and navigation method for mobile robots, using the Djikstra and A* algorithms to perform navigation between nodes. Zuo et al. [105] introduced the Voronoi diagram into the topological mapping model. On this basis, a three-stage fast path search and navigation method was designed. Topological maps allow UGVs to perform efficient path planning and navigation. However, due to the large amount of debris in fire and rescue scenarios, it is difficult for topological maps to accurately draw the obstacles and feasible paths, making it challenging for UGVs to safely navigate based on topological maps.
@@ V-B A single UAV with multi-UGVs
The collaboration of a UAV with multi-UGVs can improve the efficiency of mission execution and provide faster response capabilities. In this configuration, the UAV can provide comprehensive environmental information to UGVs, allowing UGVs to collaboratively perform search and rescue missions [132]. In addition, UGVs can also share map data, sensor information, etc. to improve collaborative work efficiency.
Miller et al. [45] proposed a semantic-based air-ground collaborative robot system framework that can realize collaborative navigation of a UAV and multiple UGVs. The system deploys a UAV to create a semantic map of the ground in real time, which is then used for UGVs’ positioning and semantic navigation tasks. To improve the navigation efficiency of multi-UGVs missions, Zhang et al. [133] suggested a strategy to organize multi-UGVs formations. In this system, a frontier-based exploration strategy is used to search for cooperative targets on the map, and the A* algorithm is then employed to plan the collision-free path through all targets. Similarly, Zang et al. [134] designed a multi-UGVs system solution based on coordinated behavior planning and trajectory planning. However, the coordination problem between UAV and multiple UGVs is not considered. In addition, in the fire and rescue missions, UGVs may be scattered over a large area, which may affect their collaboration due to the limited perception capabilities. In this case, UAV not only needs to perform map construction to assist UGV navigation, but also needs to act as a messenger to support indirect communication of UGV. Yulong et al. [122] addressed the problem of path planning for a messenger UAV in a system involving aerial and ground collaboration with multiple UGVs. This provides a solution for achieving one-to-many air-ground collaborative navigation, facilitating real-time information sharing and coordinated rescue operations in fire and rescue scenarios. To solve the path planning problem of collaborative heterogeneous robots consisting of a UAV and multiple UGVs, Ding et al. [123] developed an effective memetic algorithm to find the shortest route that allows the messenger UAV to access all moving UGVs. This treatment allows the UGVs to perform fire and rescue missions in a collective manner over a large area (refer to Fig. 14), and improve mission efficiency to a certain extent. However, for large-scale and complex fire and rescue scenarios, this kind of a single UAV with multi-UGVs collaboration requires UAV to have long-term perception and environmental modeling capabilities, which in turn makes the single UAV face challenges such as long flight time.
@@ V-C Multi-UAVs with a single UGV
The configuration of multi-UAVs with a single UGV is suitable for fire and rescue missions that require large-scale monitoring and data collection. Multiple UAVs work together to provide a more comprehensive perspective and perception capabilities, which can complete map construction faster to support UGV navigation and task execution. In addition, multiple UAVs also increase the fault tolerance and robustness of the collaborative system to a certain extent.
Stampa et al. [124] designed an air-ground collaborative system consisting of multiple UAVs and a UGV as a mobile base station, as depicted in Fig. 15. In this system, the data collected by the UAVs are fused together, and then a 3-D map of the environment is constructed based on SLAM technology to assist the navigation of robots and human rescue missions. When the accurate environmental map is completed, the system switches to monitoring mode, with some UAVs providing a real-time aerial overview of the scene and quickly detecting environmental changes to effectively address the complexity and dynamics of the fire and rescue environment. Scaramuzza et al. [135] described a collaborative system consisting of multiple UAVs that can achieve autonomous navigation in large, unknown and GPS-free environments. These works mainly involve map construction of multi-UAV systems, but pay less attention to UGV navigation based on maps. For the path planning problem in multi-UAVs-UGV collaborative mode, Ren et al. [125] developed a framework for path planning in an air-ground collaborative system consisting of a pair of UAVs and a UGV. This architecture can be used for tasks such as collaborative search and rescue of heterogeneous robotic systems. Regarding the collaborative operation of UGV-UAV, Peng et al. [126] gave a routing and scheduling algorithm for multiple UAVs and a UGV, which allows multiple UAVs carried by the UGV to perform rescue tasks such as delivering multiple packages in different locations at the same time. The above work supports the collaborative execution of firefighting and rescue missions by UAVs and a UGV in large-scale scenarios. Although multiple UAVs can provide efficient environmental perception and map construction, the efficiency and timeliness of a single UGV in executing such emergency tasks cannot be guaranteed.